OpenStudy (anonymous):

A clock radio is rated as 30 W of power output. If the radio also draws 30 W at 120 V, which will the current draw be? 0.25 A 4 A 3,600 A 0.5 A

OpenStudy (anonymous):

@andreacastillol @Adi3

OpenStudy (anonymous):

i dont know this answer

OpenStudy (michele_laino):

the electrical power is related to voltage and current through this formula: \[\Large W = V \times I\]

OpenStudy (michele_laino):

so, we can write: \[\Large I = \frac{W}{V} = \frac{{30}}{{120}} = ...?Amperes\]

OpenStudy (anonymous):

it's .25

OpenStudy (anonymous):

why am I tagged?